OpenStudy (anonymous):

1.Find k such that the line passing through (2, 1/3) and (k, 1) is parallel to the line y = 5 – 2x.

OpenStudy (campbell_st):

well the slope of the given line is m = -2 and the 2 given points have a slope of -2 so you need \[-2 = \frac{1 - \frac{1}{3}}{k - 2}\] solve for k.

OpenStudy (anonymous):

im guessing you have to do algebra? so i got -2(1-1/2)+2=k Also how did you get that?